EXS Manufacturing has agreed to pay CEO Joe Marcellicompensation in 2004 of $200,000 plus 15% of company profits. If EXS has profits of $6 million in 2004 and $4 million in 2003, how much will Joe receive in 2004? Provide justification that this amount is reasonable by first calculating how much he would receive with 2003 profit.
Answer #1

Compensation = 200,000 + 15% of company profits

Compensation in 2003 = 200,000 + (15%*4,000,000) = 800,000

Compensation in 2003 = 200,000 + (15%*6,000,000) = 1,100,000

Company profit is going up from 4 million to 6 million, an increase of 50% whereas compensation is increasing by 0.8 million to 1.1 million, an increase of 37.50%, so the compensation amount seems reasonable.